OpenStudy (anonymous):

The educational institute has a list of examiners of 150 person out of these 50 are women .125 of them know hindi and the remaining do not know. 90 of them are working and remaining are retired .what is the probability of selecting a hindi knowing working women teacher?

OpenStudy (anonymous):

50 out of 415 if i am not mistaken

OpenStudy (math&ing001):

The probability of selecting a hindi knowing working women teacher is: 50/150 * 125/150 * 90/150 = 562500/3375000 = 1/6

OpenStudy (anonymous):

nice work....big numbers lol